Circular No. 8425

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2004fe, 2004ff, 2004fg Further to IAUC 8422, H. Pugh, S. Park, and W. Li report the LOSS

discoveries of three apparent supernovae on unfiltered KAIT images:

SN  2004 UT  α_{2000}  δ_{2000}  Mag.  Offset
2004fe Oct. 30.29 0^{h}30^{m}11.^{s}27 + 2^{o}05′23′′.5 18.1 8′′.7 E, 12′′.3 S
2004ff Oct. 30.40 4 58 46.19 −21 34 12.0 18.0 11′′.2 W, 2′′.8 S
2004fg Oct. 31.19 23 47 24.45 +28 23 22.4 18.2 0′′.1 E, 14′′.7 S

Additional photometry: SN 2004fe in NGC 132, Oct. 22.28 UT, [19.0; 31.30, 17.7. SN 2004ff in ESO 552-G40, Oct. 13.41, [19.0; 21.40, [18.0; 28.39, 17.8; 31.41, 18.0. SN 2004fg in MCG +05-56-7, Oct. 23.20, [19.5; Nov. 1.17, 18.3.

SUPERNOVA 2004ew IN ESO 153-G17 M. Salvo, Australian National University (ANU); A. Pastorello, Arcetri

Observatory; and B. Schmidt, ANU, report that a low-S/N spectrum (range 400–900 nm) of SN 2004ew (cf. IAUC 8418), obtained with the ANU 2.3-m telescope (+ Double-Beam Spectrograph) at Siding Spring on Oct. 22.64 UT, shows this object to be a type-Ib (or possibly IIb) supernova, resembling SN 2000H (cf. IAUC 7375) a few weeks past maximum light. The spectrum shows the prominent He I lines at 447.2, 587.6, 667.8, 706.5, and 728.1 nm with velocities of ∼ 8200 km/s. A relatively faint, high-velocity Hα absorption is also visible (v ∼ 13300 km/s).

COMET C/2004 T3 (SIDING SPRING) Improved parabolic orbital elements from MPEC 2004-U39:

T = 2004 Oct. 29.670 TT  ω = 288.^{o}812
Ω = 48.116 }2000.0
q = 9.54814 AU  i = 72.693

COMET C/2004 Q1 (TUCKER) Visual total-magnitude estimates: Sept. 24.85 UT, 11.5 (A. Pearce,

Noble Falls, W. Australia, 0.20-m reflector); Oct. 4.88, 11.5 (A. Kammerer, Malsch, Germany, 0.30-m reflector); 10.27, 10.7 (C. S. Morris, Fillmore, CA, 0.25-m reflector); 15.52, 10.2 (S. Yoshida, Gunma, Japan, 0.40-m reflector); 23.86, 9.9 (E. P. Bus, Moncarapacho, Portugal, 10×56 binoculars).

2004 November 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green